40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jul 8 19:31:53 2012, yuta, Update, Locking, characterizing LSC arm lock by ALS error signal XarmLengthspectra20120708.pngLSCPOYarmIRlockOLTF.pngYarmLengthspectra20120708.pngLSCPOXarmIRlockOLTF.png
    Reply  Mon Jul 9 05:02:58 2012, yuta, Update, Locking, adjusted ALS filters, current RMS ALSXarmLengthspectra20120708.pngALSXarmOLTF.pngALSYarmOLTF.pngALSYarmLengthspectra20120708.png
Message ID: 6941     Entry time: Mon Jul 9 05:02:58 2012     In reply to: 6940
Author: yuta 
Type: Update 
Category: Locking 
Subject: adjusted ALS filters, current RMS 

I adjusted filters of ALS to give more phase margin.
RMS of stabilized X/Y arm length is 97 pm and 65 pm respectively.

X arm ALS:
- Openloop transfer function
UGF ~160 Hz, phase margin 30 deg
1600 usec delay (LSC-XARM had 1800 usec delay)     500 usec delay (LSC-XARM had 570 usec delay) - Edited by Yuta on July 9

ALSXarmOLTF.png

- Arm length spectra
   Red is the free run spectrum. Measured using C1:ALS-BEATX_FINE_PHASE_OUT. Calibration factor is 1.32 nm/deg.
   Green is the out-of-loop spectrum. Measured using C1:LSC-POX11_I_ERR. Calibration factor is 3.8e12 counts/m.
   Blue is the in-loop spectrum. Measured using C1:ALS-BEATX_FINE_PHASE_OUT.
   Black is the expected spectrum from openloop transfer function, such as (free run)/|1+G|.
ALSXarmLengthspectra20120708.png


   Out-of-loop estimation of RMS during X ALS is 97 pm.
   RMS mostly comes from 1 Hz and 3.3 Hz peak.
   Out-of-loop and in-loop agrees at around 10-20 Hz.

Y arm ALS:
- Openloop transfer function
UGF ~130 Hz, phase margin 20 deg
2400 usec delay (LSC-XARM had 1800 usec delay)     760 usec delay (LSC-XARM had 570 usec delay) - Edited by Yuta on July 9

ALSYarmOLTF.png

- Arm length spectra
   Red is the free run spectrum. Measured using C1:ALS-BEATY_FINE_PHASE_OUT. Calibration factor is 1.30 nm/deg.
   Green is the out-of-loop spectrum. Measured using C1:LSC-POY11_I_ERR. Calibration factor is 1.4e12 counts/m.
   Blue is the in-loop spectrum. Measured using C1:ALS-BEATY_FINE_PHASE_OUT.
   Black is the expected spectrum from openloop transferfunction, such as (free run)/|1+G|.
ALSYarmLengthspectra20120708.png

   Out-of-loop estimation of RMS during X ALS is 65 pm.
   RMS mostly comes from 1 Hz and 3.3 Hz peak.
   Out-of-loop and in-loop agrees at around 3-40 Hz.

ELOG V3.1.3-